OpenStudy (anonymous):

The polynomial 0.0013x^3-1.7x+12 models natural gas consumption in trillion cubic ft. where x=0 corresponds to 1960, x=1 to 1961, and so on. In which year between 1970 and 1990 was natural gas consumption about 17.3 trillion cubic ft.

OpenStudy (anonymous):

based on what you have written, you need to solve 0.0013x^3-1.7x+12 = 17.3

OpenStudy (anonymous):

0.0013x^3-1.7x - 5.3 = 0 13x^3 - 17000x - 53000 = 0

OpenStudy (anonymous):

perhaps you can use your calculator to solve this cubic equation